OpenStudy (anonymous):

Find the value of x in this proportion. 9/2 9/8 27/4 18

OpenStudy (anonymous):

which proportion ?>??

OpenStudy (yanasidlinskiy):

\(\cal\Huge\color{lime}{Welcome~to~Openstudy!}\)

OpenStudy (anonymous):

\[\frac{ x+3 }{ 6 }=\frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

(x+3)4=5(6) 4x+12=30 4x=18 x=9/2